# CoT-based Synthesizer: Enhancing LLM Performance through Answer Synthesis

### v1

@@ Abstract
Current inference scaling methods, such as Self-consistency and Best-of-N, have proven effective in improving the accuracy of LLMs on complex reasoning tasks. However, these methods rely heavily on the quality of candidate responses and are unable to produce correct answers when all candidates are incorrect. In this paper, we propose a novel inference scaling strategy, CoT-based Synthesizer, which leverages CoT reasoning to synthesize superior answers by analyzing complementary information from multiple candidate responses, even when all candidate responses are flawed. To enable a lightweight and cost-effective implementation, we introduce an automated data generation pipeline that creates diverse training data. This allows smaller LLMs trained on this data to improve the inference accuracy of larger models, including API-based LLMs. Experimental results across four benchmark datasets with seven policy models demonstrate that our method significantly enhances performance, with gains of 11.8% for Llama3-8B and 10.3% for GPT-4o on the MATH dataset. The corresponding training data and code are publicly available on the repository [^1] https://github.com/RUCKBReasoning/CoT-based-Synthesizer.
@@ 1 Introduction
A common inference scaling strategy is Best-of-N (BoN), which involves training a verifier to score multiple candidate responses generated by the model and selecting the highest-scoring answer as the final result Stiennon et al. (2020); Cobbe et al. (2021); Lightman et al. (2024). However, the scoring process for each candidate response in Best-of-N is independent, which fails to leverage the potential relationships between the candidates. This may lead to insufficient information integration and a susceptibility to reward hacking Skalse et al. (2022); Singhal et al. (2023). Since candidates are evaluated in isolation, the lack of cross-sample comparison makes it hard to detect outputs that exploit flaws in the reward function Christiano et al. (2017); Ouyang et al. (2022). Another widely used inference scaling strategy is Self-consistency (SC), which takes a distinct approach by selecting the most frequently generated answer from multiple inferences as the final result Wang et al. (2022). Unlike BoN, SC does not require an additional scoring step by a verifier. Instead, it uses a majority-voting process based on exact matches of candidate responses. However, this reliance on exact matching limits SC’s applicability in open-ended tasks. To address this limitation, the Universal Self-Consistency (USC) Chen et al. (2023) leverages the generative capabilities of LLMs for voting, replacing exact matching in SC, thus expanding its applicability across a wider range of tasks. However, when the correct answer appears at a low frequency among the candidates, consistency methods may fail to select the correct answer. In summary, both BoN and consistency methods operate on the assumption that the correct answer exists within the candidate set. When all candidate responses are incorrect, these methods fail to generate the correct answer. Given that BoN and USC both rely on post-processing models to evaluate candidate responses, a natural question emerges: Can we go beyond selecting from existing candidates and instead synthesize a better answer by combining the strengths of multiple candidate responses?
@@
We conduct experiments on 4 benchmarks and evaluate our method using 7 models with varying parameter scales for candidate response generation. The results demonstrate that small models trained using the data generation pipeline can not only effectively enhance the performance of candidate responses generated by various large models, even API models (e.g., with gains of $11.8\%$ for Llama3-8B and $10.3\%$ for GPT-4o on MATH), but also outperform traditional methods such as Self-consistency and Best-of-N.
Our contributions can be summarized as follows: 1) We propose a novel inference scaling strategy: CoT-based Synthesizer, which synthesizes better results by analyzing candidate responses. 2) We design an automated data generation pipeline to enable training smaller models for superior answer synthesis. 3) We validate the effectiveness of our method across multiple NLP tasks, achieving significant improvements over existing methods in complex reasoning scenarios.
@@ 2 Related Work
Self-consistency is another common inference scaling method, which operates under the assumption that the most consistent answer, determined through a voting mechanism across multiple reasoning paths, is more likely to be correct. Methods based on consistency have demonstrated significant performance improvements across various domains like mathematics Wang et al. (2022), code generation Li et al. (2022); Shi et al. (2022) and open-ended question answering Chen et al. (2023). Our method shares similarities with Chen et al. (2023) in utilizing LLMs for inference scaling, while our method is distinguished by its emphasis on the generation of novel answers. Notably, even when all candidate responses are incorrect, our method retains the capability to yield accurate outcomes.
@@
Recent studies Farinhas et al. (2023); Vernikos et al. (2023) have explored answer synthesis by combining candidate responses. Additionally, Vernikos and Popescu-Belis (2024) introduces methods that leverage quality estimation metrics to efficiently aggregate outputs from LLMs. These approaches have successfully validated the efficacy of synthesizing high-quality responses in translation tasks. Similarly, ensemble methods have been explored for combining candidate responses from different LLMs. One such method trains a ranker to conduct pairwise comparisons and selects suitable candidates to fuse into a superior answer Jiang et al. (2023). However, our work focuses on synthesizing candidate responses generated exclusively by a single LLM without using diverse models.
@@ 4 Methodology
As illustrated in Figure 2, we present an overview of our methodology. Specifically, we propose an inference scaling strategy based on the CoT reasoning, enabling the model to analyze and synthesize information across multiple candidate responses. To enable training a smaller yet effective LLM, we devise a two-stage data generation pipeline to construct a high-quality training dataset, which combines automated candidate generation with iterative repair mechanisms to improve data reliability. Finally, leveraging the generated dataset, we train Synthesizer-8B that integrates CoT reasoning to synthesize the final answer.
@@ 4.1 Synthesizer Inference
We input the generated $(x,R)$ into the CoT-based Synthesizer, which aims to analyze candidate responses, extract accurate information, and ultimately produce a high-quality response. The process involves the following steps: (1) Response Analyis: Synthesizer first conducts a thorough analysis of the relationship between the user query and each candidate response. It takes into account the frequency, relevance, and accuracy of similar responses to user queries. While high-frequency answers may indicate higher credibility, the Synthesizer does not overly rely on frequency as an evaluation criterion. Instead, it prioritizes logical coherence and factual accuracy, making it possible to identify even partially correct or misleading answers and extract valuable information from them. (2) Response Synthesis: When a correct answer exists within the candidate set, the Synthesizer further enhances and enriches the final response by incorporating effective reasoning steps from other candidate answers. In cases where there is a lack of clearly correct options or all candidate answers have deficiencies, the Synthesizer leverages its reasoning capabilities to integrate reasonable elements from multiple candidates, constructing a more coherent and accurate answer. These processes are achieved by designing specific prompts for the Synthesizer, enabling structured reasoning and accurate response generation. Detailed information can be seen in Box C.
@@ 4.2 Data Generation Pipeline
For a given training query $x$, we follow the diverse response generation process described in Section 4.1, using a sampling LLM (e.g., Llama3-8B-Instruct) to obtain the pair $(x,R)$. Then we input the pair into a high-performance response LLM (e.g., Llama3.1-70B-Instruct). Given the complexity of $x$, the response LLM may not synthesize an accurate answer on the first attempt. To address this problem, we perform multiple samplings, significantly increasing the likelihood of obtaining a correct response Tong et al. (2024). Specifically, we sample the response LLM $N$(e.g., $N=50$) times to conduct response analysis and synthesis as described in Section 4.1, generating a diverse set of synthetic answers. To ensure the quality of these synthetic answers, we filter responses using the gold answer $y_{g}$, retaining correct responses $y$, which contain CoT analysis and synthesis of $R$.
@@ 5 Experiment
Table 1: The main result of our approach and other baselines over GSM8k, MATH, WikiTQ and FeTaQA. The top two performances are highlighted in bold and underlined.
@@ 5.1 Experimental Setup
We focus on these tasks because both require substantial reasoning capabilities, and current LLMs still exhibit significant room for improvement in these domains. For the TableQA datasets, which require preprocessing to ensure format compatibility, we directly utilized the processed WikiTQ and FeTaQA test sets released by TableLLM Zhang et al. (2024b). Since our model is fine-tuned to adhere to a specific format, we extract answers that align with these format requirements from the LLM’s outputs to perform an exact match (EM) evaluation against the gold answers. Therefore, on the GSM8k, MATH, and WikiTQ datasets, we use EM to calculate accuracy. Since the gold answers in the FeTaQA dataset are in the form of complete sentences, unlike the short answer phrases in WikiTQ, we select Rouge-L as the evaluation metric to better assess the quality of generated textual answers.
@@
  • Consistency-based methods: Self-consistency (SC) Wang et al. (2022) and Universal Self-consistency (USC) Chen et al. (2023). We use SC only on GSM8k and MATH since their answers are easy to extract in complete form, allowing majority voting based on exact match. Additionally, we adopt USC across all benchmarks.
  • Best-of-N methods: We first employ the ArmoRM Wang et al. (2024) built on Llama3-8B, which is the top-ranked Llama3-8B-based model on RewardBench Lambert et al. (2024). To ensure a fairer comparison, we also train task-specific scalar reward model (Scalar RM) for MATH and WikiTQ datasets, which is the same as that used in our Synthesizer training, following the method in Cobbe et al. (2021).
  • Synthesis methods: LMCOR Vernikos et al. (2023), which directly uses the gold answers $y_{g}$ from the benchmarks as the target responses for training its synthesis model.
For our method, as described in Section 4.2, we utilize Llama3-8B-Instruct to generate candidate responses and Llama3.1-70B-Instruct to produce synthetic answers. Data generation is conducted on the training sets of MATH and WikiTQ, tailored to mathematical reasoning and TableQA. During the data filtering stage, for mathematical reasoning tasks, we filter out incorrect answers by performing an exact match between the synthesized answers and the gold answer. For TableQA, we follow TableLLM Zhang et al. (2024b), using CritiqueLLM Ke et al. (2023) to score the synthesized answers and retain only high-quality responses based on the scores, as the generated answers are often lengthy and challenging to precisely match. Subsequently, we train the CoT-based Synthesizer on Llama3-8B-Instruct Dubey et al. (2024), as it demonstrates robust general capabilities and has not yet saturated the current datasets.
Moreover, for all the aforementioned methods, we use Llama3.1-70B-Instruct as the backbone for tuning-free methods, and Llama3-8B-Instruct for training-based methods. Responses are generated by sampling k = 5 times from the policy model with a temperature of 0.9 and a Top-P value of 0.9. Further details regarding experimental settings and implementation are provided in Appendix B.
@@ 5.2 Main Results
Our method consistently outperforms baselines across benchmarks. Across the four benchmarks, Self-consistency, ArmoRM, Scalar RM and our method generally improve the CoT outputs of various policy models. We evaluate a total of seven policy models. 1) On the MATH and WikiTQ datasets, our method (including both Llama3.1-70B-Instruct and Synthesizer-8B) outperforms all baseline methods on at least six policy models, achieving an average performance that is 2% higher than the best-performing baseline. 2) Synthesizer-8B, demonstrates substantial improvements on larger models, such as performance gains of 9.6% for GLM-4-Plus and 1.6% for GPT-4o on GSM8k, even surpassing Llama3.1-70B-Instruct.
@@
Synthesizer-8B demonstrates robustness across structurally distinct models. Synthesizer-8B relies solely on Llama3-8B-Instruct to generate candidate responses and uses Llama3.1-70B-Instruct for generating replies during training. Despite this, our method still shows significant improvements across a wide range of policy models, including GLM-4-Plus, GPT-4o, Qwen2-7B-Instruct, and Qwen2.5-14B-Instruct, which are structurally and behaviorally distinct from the models used during training.
@@ 5.3 Ablation Studies
Omitting CoT training affects robustness and generalization. In contrast to the original approach, where CoT plays a critical role, we directly utilize the Llama3.1-70B-Instruct model to provide answers to each question, generating high-quality responses. These responses are subsequently used as labels in the training data, thus skipping the CoT analysis of candidate responses. Table 2 indicates that skipping CoT training slightly improves average performance compared to the CoT-prompting baseline but remains less effective than Synthesizer-8B. Notably, on some policy models, such as Llama3.1-70B-Instruct and Qwen2.5-14B-Instruct, the performance of skipping CoT training is even worse than the CoT-prompting, while on Llama3-8B-Instruct testset it surpasses the Synthesizer-8B. This suggests that removing CoT can lead to overfitting to the candidate response generation model, thereby reducing transferability and generalization across different policy models.
Omitting training on generation data impacts reasoning and synthesis capabilities. To evaluate the effectiveness of our data generation pipeline, we directly utilized an untrained Llama3-8B-Instruct as the Synthesizer to regenerate candidate responses from various policy models on mathematical reasoning tasks. Table 2 illustrates the impact of removing synthetic data. On the MATH dataset, the untrained model shows some synthesis improvements compared to CoT-prompting for some certain policy models, such as performance gains of 8% for Llama3-8B-Instruct and 3.6% for Llama3.1-8B-Instruct. However, for stronger models such as Llama3.1-70B-Instruct and Qwen2.5-14B-Instruct, the performance through answer synthesis using the untrained model is even inferior to CoT-prompting. Similarly, on the GSM8k dataset, the untrained model failed to enhance most policy models, indicating that it still makes errors during reasoning and synthesis.

### v2

@@ Abstract
Current inference scaling methods, such as Self-consistency and Best-of-N, have proven effective in improving the accuracy of LLMs on complex reasoning tasks. However, these methods rely heavily on the quality of candidate responses and are unable to produce correct answers when all candidates are incorrect. In this paper, we propose a novel inference scaling strategy, CoT-based Synthesizer, which leverages CoT reasoning to synthesize superior answers by analyzing complementary information from multiple candidate responses, even when all candidate responses are flawed. To enable a lightweight and cost-effective implementation, we introduce an automated data generation pipeline that creates diverse training data. This allows smaller LLMs trained on this data to improve the inference accuracy of larger models, including API-based LLMs. Experimental results across four benchmark datasets with seven policy models demonstrate that our method significantly enhances performance, with gains of 11.8% for Llama3-8B and 10.3% for GPT-4o on the MATH500 dataset. The corresponding training data and code are publicly available on the repository [^1] https://github.com/RUCKBReasoning/CoT-based-Synthesizer.
@@ 1 Introduction
A common inference scaling strategy is Best-of-N (BoN), which involves training a verifier to score multiple candidate responses generated by the model and selecting the highest-scoring answer as the final result Stiennon et al. (2020); Cobbe et al. (2021); Lightman et al. (2024). However, the scoring process for each candidate response in Best-of-N is independent, which fails to leverage the potential relationships between the candidates. This may lead to insufficient information integration and a susceptibility to reward hacking Skalse et al. (2022); Singhal et al. (2023). Since candidates are evaluated in isolation, the lack of cross-sample comparison makes it hard to detect outputs that exploit flaws in the reward function Christiano et al. (2017); Ouyang et al. (2022). Another widely used inference scaling strategy is Self-consistency (SC), which takes a distinct approach by selecting the most frequently generated answer from multiple inferences as the final result Wang et al. (2022). Unlike BoN, SC does not require an additional scoring step by a verifier. Instead, it uses a majority-voting process based on exact matches of candidate responses. However, this reliance on exact matching limits SC’s applicability in open-ended tasks. To address this limitation, the Universal Self-Consistency (USC) Chen et al. (2023) leverages the generative capabilities of LLMs for voting, replacing exact matching in SC, thus expanding its range of applications. However, when the correct answer appears at a low frequency among the candidates, consistency methods may fail to select the correct answer. In summary, both BoN and consistency methods operate on the assumption that the correct answer exists within the candidate set. When all candidate responses are incorrect, these methods fail to generate the correct answer. Given that BoN and USC both rely on post-processing models to evaluate candidate responses, a natural question emerges: Can we go beyond selecting from existing candidates and instead synthesize a better answer by combining the strengths of multiple candidate responses?
@@
We conduct experiments on 4 benchmarks and evaluate our method using 7 models with varying parameter scales for candidate response generation. The results demonstrate that small models trained using the data generation pipeline can not only effectively enhance the performance of candidate responses generated by various large models, even API models (e.g., with gains of $11.8\%$ for Llama3-8B and $10.3\%$ for GPT-4o on MATH500), but also outperform traditional methods such as Self-consistency and Best-of-N.
Our contributions can be summarized as follows: 1) We propose a novel inference scaling strategy: CoT-based Synthesizer, which synthesizes better results by analyzing candidate responses. 2) We design an automated data generation pipeline to enable training smaller efficient models for superior answer synthesis. 3) We validate the effectiveness of our method across multiple NLP tasks, achieving significant improvements over existing methods in complex reasoning scenarios.
@@ 2 Related Work
Self-consistency is another common inference scaling method, which operates under the assumption that the most consistent answer, determined through a voting process, is more likely to be correct. Methods based on consistency have demonstrated significant performance improvements across various domains like mathematics Wang et al. (2022), code generation Shi et al. (2022); Li et al. (2022) and open-ended question answering Chen et al. (2023). Our method shares similarities with Chen et al. (2023) in utilizing LLMs for inference scaling, while our method is distinguished by its emphasis on the generation of novel answers. Notably, even when all candidate responses are incorrect, our method retains the capability to yield accurate outcomes.
@@
Recent studies Farinhas et al. (2023); Vernikos et al. (2023) have explored answer synthesis by combining candidate responses. Additionally, Vernikos and Popescu-Belis (2024) proposes methods that employ quality estimation metrics to effectively combine outputs from LLMs. These approaches have successfully validated the efficacy of synthesizing high-quality responses in translation tasks. Similarly, ensemble methods have been explored for combining candidate responses from different LLMs. One such method trains a ranker to conduct pairwise comparisons and selects suitable candidates to fuse into a superior answer Jiang et al. (2023). However, our work focuses on synthesizing candidate responses generated exclusively by a single LLM without using diverse models.
@@ 4 Methodology
As illustrated in Figure 2, we present an overview of our methodology. Specifically, we propose an inference scaling strategy based on the CoT reasoning, enabling the model to analyze and synthesize information across multiple candidate responses. To enable training a smaller yet effective LLM, we devise a two-stage data generation pipeline to construct a high-quality training dataset, which combines automated candidate generation with an extra repair step to improve data reliability. Finally, leveraging the generated dataset, we train Synthesizer-8B that integrates CoT reasoning to synthesize the final answer. Furthermore, our approach incorporates various LLMs, including the sampling LLM, response LLM, policy model, and base model, each playing a distinct role in inference and training. Formal definitions of these components are provided in Appendix D.
@@ 4.1 Synthesizer Inference
We input the generated $(x,R)$ into the CoT-based Synthesizer, which aims to analyze candidate responses, extract accurate information, and ultimately produce a high-quality response. The process involves the following steps: (1) Response Analyis: Synthesizer first conducts a thorough analysis of the relationship between the user query and each candidate response. It takes into account the frequency, relevance, and accuracy of similar responses to user queries. While high-frequency answers may indicate higher credibility, the Synthesizer does not overly rely on frequency as an evaluation criterion. Instead, it prioritizes logical coherence and factual accuracy, making it possible to identify even partially correct or misleading answers and extract valuable information from them. (2) Response Synthesis: When a correct answer exists within the candidate set, the Synthesizer further enhances and enriches the final response by incorporating effective reasoning steps from other candidate answers. In cases where there is a lack of clearly correct options or all candidate answers have deficiencies, the Synthesizer leverages its reasoning capabilities to integrate reasonable elements from multiple candidates, constructing a more coherent and accurate answer. These processes are achieved by designing specific prompts for the Synthesizer, enabling structured reasoning and accurate response generation. Detailed information can be seen in Box E.
@@ 4.2 Data Generation Pipeline
For a given training query $x$, we follow the diverse response generation process described in Section 4.1, using a sampling LLM (e.g., Llama3-8B-Instruct) to obtain the pair $(x,R)$. Then we input the pair into a high-performance response LLM (e.g., Llama3.1-70B-Instruct), which employs a CoT reasoning process to generate responses. Given the complexity of $x$, the response LLM may not synthesize an accurate answer on the first attempt. To address this problem, we perform multiple samplings, significantly increasing the likelihood of obtaining a correct response Tong et al. (2024). Specifically, we sample the response LLM $N$(e.g., $N=50$) times to conduct response analysis and synthesis as described in Section 4.1, generating a diverse set of synthetic answers. To ensure the quality of these synthetic answers, we filter responses using the gold answer $y_{g}$, retaining correct responses $y$, which contain CoT analysis and synthesis of $R$.
@@ 5 Experiment
Table 1: The main result of our approach and other baselines over GSM8k, MATH500, WikiTQ and FeTaQA. The top two performances are highlighted in bold and underlined.
@@ 5.1 Experimental Setup
We focus on these tasks because both require substantial reasoning capabilities, and current LLMs still exhibit significant room for improvement in these domains. For the TableQA datasets, which require preprocessing to ensure format compatibility, we directly utilized the processed WikiTQ and FeTaQA test sets released by TableLLM Zhang et al. (2024b). Since our model is fine-tuned to adhere to a specific format, we extract answers that align with these format requirements from the LLM’s outputs to perform an exact match (EM) evaluation against the gold answers. Therefore, on the GSM8k, MATH500, and WikiTQ datasets, we use EM to calculate accuracy. Since the gold answers in the FeTaQA dataset are in the form of complete sentences, unlike the short answer phrases in WikiTQ, we select Rouge-L as the evaluation metric to better assess the quality of textual answers.
@@
  • Consistency-based methods: Self-consistency (SC) Wang et al. (2022) and Universal Self-consistency (USC) Chen et al. (2023). We use SC only on GSM8k and MATH500 since their answers are easy to extract, allowing majority voting based on exact match. Additionally, we adopt USC across all benchmarks.
  • Best-of-N methods: We first employ the ArmoRM Wang et al. (2024) built on Llama3-8B, which is the top-ranked Llama3-8B-based model on RewardBench Lambert et al. (2024). To ensure a fairer comparison, we also train task-specific scalar reward model (Scalar RM) for MATH500 and WikiTQ datasets, which is the same as that used in our Synthesizer training, following the method in Cobbe et al. (2021).
  • Synthesis methods: LMCOR Vernikos et al. (2023), which directly uses the gold answers $y_{g}$ from the benchmarks as the target responses for training its synthesis model, without incorporating candidate response analysis and integration through CoT process.
As described in Section 4.2, we utilize Llama3-8B-Instruct as the Sampling LLM to generate candidate responses and Llama3.1-70B-Instruct as the Response LLM to generate synthetic answers. Our data generation pipeline expands the original 12k MATH training samples to 295k and the 18k WikiTQ training samples (from TableLLM) to 87k. During data filtering, for mathematical reasoning tasks, we filter out incorrect answers via exact matching with the gold answer. For TableQA, we follow TableLLM, using CritiqueLLM Ke et al. (2023) to score the synthesized answers and retain only high-quality responses, as the generated answers are often lengthy and challenging to precisely match. Subsequently, we train the CoT-based Synthesizer on Llama3-8B-Instruct, as it demonstrates robust general capabilities and has not yet saturated the current datasets.
For all methods, we use Llama3.1-70B-Instruct as the backbone for tuning-free methods, and Llama3-8B-Instruct for training-based methods. Responses are generated by sampling k = 5 times from the policy model with T = 0.9 and Top-P = 0.9. To ensure reliability, each experiment is conducted three times, and the reported results are the average of these runs. Further details regarding experimental settings and implementation are provided in Appendix C.
@@ 5.2 Main Results
Our method consistently outperforms baselines across benchmarks. Across the four benchmarks, SC, ArmoRM, Scalar RM and our method generally improve the CoT outputs of various policy models. We evaluate a total of seven policy models. 1) On the MATH500 and WikiTQ datasets, our method (including both Llama3.1-70B-Instruct and Synthesizer-8B) outperforms all baseline methods on at least six policy models, achieving an average performance that is 2% higher than the best-performing baseline. 2) Synthesizer-8B demonstrates substantial improvements on larger models(e.g., GLM-4-Plus and Llama3.1-70B), outperforming all methods, even surpassing directly using Llama3.1-70B as Synthesizer.
@@
Synthesizer-8B demonstrates robustness across structurally distinct models. Synthesizer-8B relies solely on Llama3-8B-Instruct to generate candidate responses and uses Llama3.1-70B-Instruct for generating final answers. Despite this, our method still shows significant improvements across a wide range of policy models, including GLM-4-Plus, GPT-4o, Qwen2-7B, and Qwen2.5-14B, which are structurally and behaviorally distinct from the models used during training.
@@ 5.3 Ablation Studies
Omitting CoT training affects robustness and generalization. In contrast to the original approach, where CoT plays a critical role, we directly utilize the Llama3.1-70B as the Synthesizer to provide answers to each question, generating high-quality responses. These responses are subsequently used as labels in the training data, thus skipping the CoT analysis of candidate responses. Table 2 indicates that skipping CoT training slightly improves average performance over CoT-prompting but remains inferior to Synthesizer-8B. Notably, for models like Llama3.1-70B and Qwen2.5-14B, the performance of skipping CoT training is even worse than the CoT-prompting, while on Llama3-8B testset it surpasses the Synthesizer-8B. This suggests that removing CoT can lead to overfitting to the Sampling LLM, thereby reducing generalization across different policy models.
Omitting training on generation data impacts reasoning and synthesis capabilities. To evaluate the effectiveness of our data generation pipeline, we directly utilized an untrained Llama3-8B-Instruct to regenerate candidate responses from various policy models on mathematical reasoning tasks. Table 2 illustrates that while the untrained model shows some synthesis improvements compared to CoT-prompting on MATH500 for some models (e.g., +8% on Llama3-8B, +3.6% on Llama3.1-8B), it underperforms for stronger models like Llama3.1-70B and Qwen2.5-14B. Similarly, on the GSM8k dataset, Llama3-8B fails to enhance most policy models, indicating that it still makes errors during reasoning and synthesis.